solve the equation of 6x + 8 =20
Found 2 solutions by mathmandotcom, Alan3354:
Answer by mathmandotcom(31)

6x+8=20
-8 -8
6x=12
6x/6=12/6
x=6/3=2
thanks al

Answer by Alan3354(69443)

solve the equation of 6x + 8 =20
6x + 8 = 20
Subtract 8
6x = 12
Divide by 6
x = 2